\begin{thebibliography}{99}

\bibitem{AHK} E.\:Akin, M.\:Hurley, J.\:Kennedy, Dynamics of topologically generic homeomorphisms. Mem. Amer. Math. Soc. 164 (2003).

\bibitem{BGO} A.D.\:Barwell, C.\:Good, P.\:Oprocha, Shadowing and expansivity in subspaces. Fund. Math. 219 (2012), 223--243.

\bibitem{C} C.\:Conley, Isolated invariant sets and the Morse index. CBMS Regional Conference Series in Mathematics, 38. American Mathematical Society, Providence, R.I., 1978.

\bibitem{KH} A.\:Katok, B.\:Hasselblatt, 
Introduction to the modern theory of dynamical systems. Encyclopedia of Mathematics and Its Applications, 54. Cambridge Univ. Press, 1995.

\bibitem{K1} N.\:Kawaguchi, Generic and dense distributional chaos with shadowing. J. Difference Equ. Appl. 27 (2021), 1456--1481.

\bibitem{K2} N.\:Kawaguchi, Contractive shadowing property of dynamical systems. Qual. Theory Dyn. Syst. 23 (2024), 221.

\bibitem{M} T.K.S.\:Moothathu, Implications of pseudo-orbit tracing property for continuous maps on compacta. Topology Appl. 158 (2011), 2232--2239.

\bibitem{RW} D.\:Richeson, J.\:Wiseman, Chain recurrence rates and topological entropy. Topology Appl. 156 (2008), 251--261.

\bibitem{W} P.\:Walters, An introduction to ergodic theory. Graduate Texts in Mathematics, 79. Springer, 1982.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage[margin=30mm]{geometry}
\usepackage{amsmath,amssymb}
\usepackage{amsthm}
\usepackage{mathrsfs}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}{Lemma}[section]

\newtheorem{defi}{Definition}[section]
\newtheorem{ex}{Example}[section]
\newtheorem{rem}{Remark}[section]

\newtheorem*{theorem}{\it{Theorem}}

\usepackage{enumitem}

\begin{document}

\title{On ball expanding maps}
\author{Noriaki Kawaguchi}
\subjclass[2020]{37B05, 37B65}
\keywords{ball expanding, shadowing, chain recurrent set, entropy, locally eventually onto}
\address{Department of Mathematical and Computing Science, School of Computing, Institute of Science Tokyo, 2-12-1 Ookayama, Meguro-ku, Tokyo 152-8552, Japan}
\email{gknoriaki@gmail.com}

\begin{abstract}
We study the dynamical properties of ball expanding maps, a class of continuous self-maps defined on compact metric spaces. For a ball expanding map, we show that: (1) the set of periodic points is dense in the chain recurrent set; (2) if the topological entropy of the map is zero, then the chain recurrent set is finite; (3) the map has only finitely many chain components; (4) if the space is perfect, then the topological entropy of the map is positive; and (5) if the space is connected, then the map is locally eventually onto and hence mixing. Several examples are also provided.
\end{abstract}

\maketitle

\markboth{NORIAKI KAWAGUCHI}{On ball expanding maps}

\section{Main results and their proofs}

{\em Expanding maps} have played a significant role in the study of dynamical systems due to their rich structure and chaotic properties (see, e.g., \cite{KH}). {\em Ball expanding maps} are a class of continuous self-maps defined on compact metric spaces, characterized by their local expansive behavior \cite{BGO}. In this paper, we investigate the dynamics of ball expanding maps, focusing on periodic points, chain recurrence, entropy, and mixing properties.  Our results reveal how local ball expansion imposes global dynamical constraints. Specifically, for a ball expanding map, we show that
\begin{itemize}
\item[--] the set of periodic points is dense in the chain recurrent set (Theorem 1.1),
\item[--] if the topological entropy of the map is zero, then the chain recurrent set is finite (1.2),
\item[--] the map has only finitely many chain components (1.3),
\item[--] if the space is perfect, then the topological entropy of the map is positive (1.4),
\item[--] if the space is connected, then the map is locally eventually onto and hence mixing (1.5).
\end{itemize}

We begin with the definition of ball expanding maps. Throughout, $X$ denotes a compact metric space endowed with a metric $d$. For $x\in X$ and $r>0$, we denote by $B_r(x)$ the closed $r$-ball centered at $x$: $B_r(x)=\{y\in X\colon d(x,y)\le r\}$.

\begin{defi}
\normalfont
A continuous map $f\colon X\to X$ is said to be {\em ball expanding} if there are $0<L<1$ and $\delta_0>0$ such that
\[
B_\delta(f(x))\subset f(B_{L\delta}(x))
\]
for all $x\in X$ and $0<\delta\le\delta_0$ \cite{BGO}.
\end{defi}

\begin{rem}
\normalfont
\begin{itemize}
\item By Theorem 4.3 of \cite{BGO}, we know that if a continuous map $f\colon X\to X$ is ball expanding, then $f$ has the {\em h-shadowing property}. A map $f\colon X\to X$ is called an {\em open map} if $f(U)$ is open for all open subset $U$ of $X$. We know that a continuous map $f\colon X\to X$ with the h-shadowing property is an open map.
\item In \cite{BGO}, a continuous map $f\colon X\to X$ is said to be {\em expanding} if there are $\delta_0>0$ and $0<L<1$ such that for any $x,y\in X$, $d(x,y)\le\delta_0$ implies
\[
d(f(x),f(y))\ge L^{-1}d(x,y).
\]
By Theorem 2.5 of \cite{BGO}, we know that for a continuous map $f\colon X\to X$, the following conditions are equivalent
\begin{itemize}
\item $f$ is expanding and open,
\item $f$ is ball expanding and locally one-to-one.
\end{itemize}
\end{itemize}

\end{rem}

In \cite{K2}, the ball expanding condition has been related to the contractive shadowing property, i.e., the Lipschitz shadowing property such that the Lipschitz constant is less than $1$. We recall the definition of shadowing.

\begin{defi}
\normalfont
Let $f\colon X\to X$ be a continuous map and let $\xi=(x_i)_{i\ge0}$ be a sequence of points in $X$. For $\delta>0$, $\xi$ is called a {\em $\delta$-pseudo orbit} of $f$ if
\[
\sup_{i\ge0}d(f(x_i),x_{i+1})\le\delta.
\]
For $\epsilon>0$, $\xi$ is said to be {\em $\epsilon$-shadowed} by $x\in X$ if
\[
\sup_{i\ge0}d(f^i(x),x_i)\le\epsilon.
\]
The map $f$ is said to have the {\em shadowing property} if for any $\epsilon>0$, there is $\delta>0$ such that every $\delta$-pseudo orbit of $f$ is $\epsilon$-shadowed by some point of $X$. For $L>0$, we say that $f$ has the {\em $L$-Lipschitz shadowing property} if there is $\delta_0>0$ such that for any $0<\delta\le\delta_0$, every $\delta$-pseudo orbit of $f$ is $L\delta$-shadowed by some point of $X$. We say that $f$ has the {\em contractive shadowing property} if $f$ has the $L$-Lipschitz shadowing property for some $0<L<1$ \cite{K2}.
\end{defi}

\begin{rem}
\normalfont
For a continuous map $f\colon X\to X$, let $0<L<1$ and $\delta_0>0$ be as in the definition of ball expanding. By Remark 1.3 (2) of \cite{K2}, we know that $f$ has the $M$-Lipschitz shadowing property, where $M=L/(1-L)$. Note that for any $i\ge0$, we have 
\[
B_\delta(f^i(x))\subset f^i(B_{L^i\delta}(x))
\]
for all $x\in X$ and $0<\delta\le\delta_0$. As a consequence, for every $i\ge1$, $f^i$ has the $M_i$-Lipschitz shadowing property, where $M_i=L^i/(1-L^i)$.
\end{rem}

{\em Chain components}, which appear in the (so-called) fundamental theorem of dynamical systems by Conley, are essential objects for a global understanding of dynamical systems \cite{C}. Our main results and their proofs are based on the notions of chain recurrence and chain components. Their definitions are as follows.

\begin{defi}
\normalfont
Let $f\colon X\to X$ be a continuous map. For $\delta>0$, a finite sequence $(x_i)_{i=0}^k$ of points in $X$, where $k\ge1$, is called a {\em $\delta$-chain} of $f$ if
\[
\sup_{0\le i\le k-1}d(f(x_i),x_{i+1})\le\delta.
\]
For any $x,y\in X$, the notation $x\rightarrow y$ means that for every $\delta>0$,  there is a $\delta$-chain $(x_i)_{i=0}^k$ of $f$ with $x_0=x$ and $x_k=y$. The {\em chain recurrent set} $CR(f)$ for $f$ is defined by
\[
CR(f)=\{x\in X\colon x\rightarrow x\}.
\]
We define a relation $\leftrightarrow$ in
\[
CR(f)^2=CR(f)\times CR(f)
\]
by: for any $x,y\in CR(f)$, $x\leftrightarrow y$ if and only if $x\rightarrow y$ and $y\rightarrow x$. Note that $\leftrightarrow$ is a closed equivalence relation in $CR(f)^2$ and satisfies $x\leftrightarrow f(x)$ for all $x\in CR(f)$. An equivalence class $C$ of $\leftrightarrow$ is called a {\em chain component} for $f$. We denote by $\mathcal{C}(f)$ the set of chain components for $f$.
\end{defi}

The basic properties of chain components are given in the following remark. A subset $S$ of $X$ is said to be {\em $f$-invariant} if $f(S)\subset S$.

\begin{rem}
\normalfont
The following properties hold
\begin{itemize}
\item $CR(f)=\bigsqcup_{C\in\mathcal{C}(f)}C$, a disjoint union,
\item every $C\in\mathcal{C}(f)$ is a closed $f$-invariant subset of $CR(f)$,
\item for all $C\in\mathcal{C}(f)$, $f|_C\colon C\to C$ is chain transitive, i.e., for any $x,y\in C$ and $\delta>0$, there is a $\delta$-chain $(x_i)_{i=0}^k$ of $f|_C$ with $x_0=x$ and $x_k=y$.
\end{itemize}
\end{rem}

\begin{rem}
\normalfont
Let $f\colon X\to X$ be a continuous map. 
\begin{itemize}
\item[(1)] It holds that $CR(f^i)=CR(f)$ for all $i\ge1$. 
\item[(2)] The following conditions are equivalent
\begin{itemize}
\item $f\colon CR(f)\to CR(f)$ is a homeomorphism,
\item $f^i\colon CR(f^i)\to CR(f^i)$ is a homeomorphism for all $i\ge1$,
\item $f^i\colon CR(f^i)\to CR(f^i)$ is a homeomorphism for some $i\ge1$.
\end{itemize}
\item[(3)] The following conditions are equivalent
\begin{itemize}
\item $\mathcal{C}(f)$ is a finite set,
\item $\mathcal{C}(f^i)$ is a finite set for all $i\ge1$,
\item $\mathcal{C}(f^i)$ is a finite set for some $i\ge1$.
\end{itemize}
\end{itemize}

(3) can be proved as follows. If $\mathcal{C}(f)$ is a finite set, then for all $i\ge1$, by $|\mathcal{C}(f^i)|\le i\cdot|\mathcal{C}(f)|$, $\mathcal{C}(f^i)$ is a finite set.  If $\mathcal{C}(f^i)$ is a finite set for some $i\ge1$, then by $|\mathcal{C}(f)|\le|\mathcal{C}(f^i)|$, $\mathcal{C}(f)$ is a finite set.
\end{rem}

Given a continuous map $f\colon X\to X$, we denote by $Per(f)$ the set of periodic points for $f$:
\[
Per(f)=\bigcup_{j\ge1}\{x\in X\colon f^j(x)=x\}.
\]
By Theorem 1.2 of \cite{K2}, we know that if $f$ has the $L$-Lipschitz shadowing property for some $0<L<1/2$, then $f$ satisfies $CR(f)=\overline{Per(f)}$. Let $0<L<1$ and $\delta_0>0$ be as in the definition of ball expanding.  By Remark 1.2, taking $i\ge1$ with $L^i<1/3$, we have that $f^i$ has the $M_i$-Lipschitz shadowing property, where $M_i=L^i/(1-L^i)<1/2$. It follows that
\[
CR(f^i)=\overline{Per(f^i)}.
\]
By $CR(f^i)=CR(f)$ and $Per(f^i)=Per(f)$, we obtain $CR(f)=\overline{Per(f)}$. In other words, we obtain the following theorem.

\begin{thm}
If a continuous map $f\colon X\to X$ is ball expanding, then $f$ satisfies $CR(f)=\overline{Per(f)}$.
\end{thm}

For a continuous map $f\colon X\to X$, we denote by $h_{\rm top}(f)$ the topological entropy of $f$ (see, e.g., \cite{W} for details). By results of \cite{M}, if $f$ has the shadowing property, then $f$ satisfies $h_{\rm top}(f)=0$ if and only if
\[
f|_{CR(f)}\colon CR(f)\to CR(f)
\]
is an equicontinuous homeomorphism. We recall some related results from \cite{K2} (see Theorems 1.4 and 1.7 of \cite{K2}).

\begin{theorem}
\normalfont
\begin{itemize}
\item If a homeomorphism $f\colon X\to X$ has the $L$-Lipschitz shadowing property for some $0<L<1$, then $X$ is a finite set.
\item For $L>0$, if a  continuous map $f\colon X\to X$ has the $L$-Lipschitz shadowing property, then so does $f|_{CR(f)}\colon CR(f)\to CR(f)$.
\end{itemize} 
\end{theorem}

From these facts, it follows that if a continuous map $f\colon X\to X$ has the $L$-Lipschitz shadowing property for some $0<L<1$, then the following conditions are equivalent
\begin{itemize}
\item $h_{\rm top}(f)=0$,
\item $CR(f)$ is a finite set,
\item $f|_{CR(f)}\colon CR(f)\to CR(f)$ is a homeomorphism.
\end{itemize}
Let $0<L<1$ and $\delta_0>0$ be as in the definition of ball expanding.  As in Remark 1.2, by taking $i\ge1$ with $L^i<1/2$, we have that $f^i$ has the $M_i$-Lipschitz shadowing property, where $M_i=L^i/(1-L^i)<1$. Note that $h_{\rm top}(f^i)=i\cdot h_{\rm top}(f)$. By the above equivalence, we obtain the following theorem (see also Remark 1.4).

\begin{thm}
If a continuous map $f\colon X\to X$ is ball expanding, then the following conditions are equivalent
\begin{itemize}
\item[(1)] $h_{\rm top}(f)=0$,
\item[(2)] $CR(f)$ is a finite set,
\item[(3)] $f|_{CR(f)}\colon CR(f)\to CR(f)$ is a homeomorphism.
\end{itemize}
\end{thm}

By Theorem 1.8 of \cite{K2}, we know that if $f$ has the $L$-Lipschitz shadowing property for some $0<L<1$, then $\mathcal{C}(f)$ is a finite set. By a similar argument as above, we obtain the following theorem (see also Remark 1.4).

\begin{thm}
If a continuous map $f\colon X\to X$ is ball expanding, then $\mathcal{C}(f)$ is a finite set.
\end{thm}

Following \cite{AHK}, we define the {\em terminal} chain components as follows.

\begin{defi}
\normalfont
Given a continuous map $f\colon X\to X$, we say that a closed $f$-invariant subset $S$ of $X$ is {\em chain stable} if for any $\epsilon>0$, there is $\delta>0$ such that every $\delta$-chain $(x_i)_{i=0}^k$ of $f$ with $x_0\in S$ satisfies $d(x_k,S)=\inf_{y\in S}d(x_k,y)\le\epsilon$. We say that $C\in\mathcal{C}(f)$ is {\em terminal} if $C$ is chain stable. We denote by $\mathcal{C}_{\rm ter}(f)$ the set of terminal chain components for $f$.
\end{defi}

\begin{rem}
\normalfont
\begin{itemize}
\item For a continuous map $f\colon X\to X$ and $x\in X$, the {\em $\omega$-limit set} $\omega(x,f)$ of $x$ for $f$ is defined as the set of $y\in X$ such that $\lim_{j\to\infty}f^{i_j}(x)=y$ for some sequence $0\le i_1<i_2<\cdots$. Note that $\omega(x,f)$ is a closed $f$-invariant subset of $X$ and satisfies
\[
\lim_{i\to\infty}d(f^i(x),\omega(x,f))=0.
\]
Since we have $y\rightarrow z$ for all $y,z\in\omega(x,f)$, there is an unique $C(x,f)\in\mathcal{C}(f)$ such that $\omega(x,f)\subset C(x,f)$ and so
\[
\lim_{i\to\infty}d(f^i(x),C(x,f))=0.
\]
\item Let $f\colon X\to X$ be a continuous map and let $S$ be a closed $f$-invariant subset of $X$. We see that $S$ is chain stable if and only if $x\rightarrow y$ implies $y\in S$ for all $x\in S$ and $y\in X$.
\item A partial order $\le$ on $\mathcal{C}(f)$ can be defined by: for any $C,D\in\mathcal{C}(f)$, $C\le D$ if only if there are $x\in C$ and $y\in D$ such that $x\rightarrow y$. We easily see that for any $C\in\mathcal{C}(f)$, $C\in\mathcal{C}_{\rm ter}(f)$ if and only if $C$ is maximal with respect to $\le$.
\end{itemize}
\end{rem}

We can show that every continuous map $f\colon X\to X$ satisfies $\mathcal{C}_{\rm ter}(f)\ne\emptyset$ (see, e.g., Lemma 2.1 and Corollary 2.1 of \cite{K1}), but this fact is fairly obvious when $\mathcal{C}(f)$ is a finite set.

In order to prove Theorems 1.4 and 1.5, let us first prove the following lemma.

\begin{lem}
Let $f\colon X\to X$ be a continuous map. Let $0<L<1$ and $\delta_0>0$ be as in the definition of ball expanding. For any $C\in\mathcal{C}(f)$ and $x\in X$, if $d(x,C)\le\delta_0$ and
\[
\lim_{i\to\infty}d(f^i(x),C)=0,
\]
then $x\in C$.
\end{lem}

\begin{proof}
Note that for any $i\ge0$, we have 
\[
B_{\delta_0}(f^i(y))\subset f^i(B_{L^i\delta_0}(y))
\]
for all $y\in X$. Since $d(x,C)\le\delta_0$, we have $d(x,z)\le\delta_0$ for some $z\in C$. We take a sequence $y_i\in C$, $i\ge0$, such that $z=f^i(y_i)$ for all $i\ge0$. For every $i\ge0$, we obtain
\[
x\in B_{\delta_0}(z)=B_{\delta_0}(f^i(y_i))\subset f^i(B_{L^i\delta_0}(y_i))
\]
and so $x=f^i(z_i)$ for some $z_i\in B_{L^i\delta_0}(y_i)$. Since $d(z_i,C)\le d(z_i,y_i)\le L^i\delta_0$ for all $i\ge0$, we have
\[
\lim_{i\to\infty}d(z_i,C)=0.
\]
It follows that $w\rightarrow x$ for some $w\in C$. On the other hand, since
\[
\lim_{i\to\infty}d(f^i(x),C)=0,
\]
we have $x\rightarrow v$ for some $v\in C$. By $v,w\in C$, we obtain $v\rightarrow w$. This with $x\rightarrow v$ implies $x\rightarrow w$. It follows that $x\rightarrow x$ and $x\leftrightarrow w$; therefore, $x\in C$. This completes the proof of the lemma.
\end{proof}

In the next remark, by using Lemma 1.1, we prove a non-trivial property of ball expanding maps.

\begin{rem}
\normalfont
Let $f\colon X\to X$ be a continuous map. Let $0<L<1$ and $\delta_0>0$ be as in the definition of ball expanding. For any $x\in X$, we take $C(x,f)\in\mathcal{C}(f)$ such that $\omega(x,f)\subset C(x,f)$ and so
\[
\lim_{i\to\infty}d(f^i(x),C(x,f))=0.
\]
Note that $d(f^i(x),C(x,f))\le\delta_0$ for some $i\ge0$. By Lemma 1.1, we obtain $f^i(x)\in C(x,f)$. Since $x\in X$ is arbitrary, we conclude that
\[
X=\bigcup_{i\ge0}f^{-i}(CR(f))
\]
(see also Example 2.1 in Section 2).
\end{rem}

By Theorem 1.3 and Lemma 1.1, we obtain the following lemma.

\begin{lem}
If a continuous map $f\colon X\to X$ is ball expanding, then every $C\in\mathcal{C}_{\rm ter}(f)$ is a clopen subset of $X$.
\end{lem}

\begin{proof}
Since $f$ is ball expanding, by Theorem 1.3, $\mathcal{C}(f)$ is a finite set. This with $C\in\mathcal{C}_{\rm ter}(f)$ implies the existence of $\delta_1>0$ such that every $x\in X$ with $d(x,C)\le\delta_1$ satisfies
\[
\lim_{i\to\infty}d(f^i(x),C)=0.
\]
Let $0<L<1$ and $\delta_0>0$ be as in the definition of ball expanding. Then, for any $x\in X$ with $d(x,C)\le\min\{\delta_0,\delta_1\}$, by Lemma 1.1, we obtain $x\in C$. This implies that $C$ is clopen in $X$, thus the lemma has been proved.
\end{proof}

The space $X$ is said to be {\em perfect} if $X$ has no isolated points, i.e., $x\in\overline{X\setminus\{x\}}$ for all $x\in X$. By Theorem 1.2 and Lemma 1.2, we obtain the following theorem.

\begin{thm}
If $X$ is perfect and if a continuous map $f\colon X\to X$ is ball expanding, then $h_{\rm top}(f)>0$.
\end{thm}

\begin{proof}
Let $C\in\mathcal{C}_{\rm ter}(f)$. Since $f$ is ball expanding, by Lemma 1.2, $C$ is a clopen subset of $X$. If $h_{\rm top}(f)=0$, then Theorem 1.2 implies that $C$ is a finite set and therefore every $x\in C$ is an isolated point of $X$. Since $X$ is perfect, we conclude that $h_{\rm top}(f)>0$, completing the proof.
\end{proof}

Let us recall here the definition of the h-shadowing property.

\begin{defi}
\normalfont
A continuous map $f\colon X\to X$ is said to have the {\em h-shadowing property} if for any $\epsilon>0$, there is $\delta>0$ such that every $\delta$-chain $(x_i)_{i=0}^k$ of $f$ satisfies
\[
\sup_{0\le i\le k-1}d(f^i(x),x_i)\le\epsilon
\]
and $f^k(x)=x_k$ for some $x\in X$.
\end{defi}

A continuous map $f\colon X\to X$ is said to be {\em mixing} if for any non-empty open subsets $U,V$ of $X$, there is $i\ge0$ such that $f^j(U)\cap V\ne\emptyset$ for all $j\ge i$. We say that $f$ is {\em locally eventually onto} if every non-empty open subset $U$ of $X$ satisfies $f^i(U)=X$ for some $i\ge0$. We easily see that if $f$ is locally eventually onto, then $f$ is mixing, and the converse holds when $f$ has the h-shadowing property.

\begin{rem}
\normalfont
Let $f\colon X\to X$ be a continuous map. We say that $f$ is {\em chain mixing} if for any $x,y\in X$ and $\delta>0$, there is $j>0$ such that for each $k\ge j$, a $\delta$-chain $(x_i)_{i=0}^k$ of $f$ satisfies $x_0=x$ and $x_k=y$. If $f$ is mixing, then $f$ is chain mixing, and the converse holds when $f$ has the shadowing property. By Corollary 14 of \cite{RW}, we know that if $X$ is connected and $f$ is chain transitive, i.e., $\mathcal{C}(f)=\{X\}$, then $f$ is chain mixing. 
\end{rem}

We shall present the final result of this section (see Appendix B for a generalization). 

\begin{thm}
If $X$ is connected and if a continuous map $f\colon X\to X$ is ball expanding, then $f$ is locally eventually onto.
\end{thm}

\begin{proof}
We take $C\in\mathcal{C}_{\rm ter}(f)$. Since $f$ is ball expanding, by Lemma 1.2, $C$ is a clopen subset of $X$. Since $X$ is connected, it follows that $X=C$. By Corollary 14 of \cite{RW}, $f$ is chain mixing. Since $f$ is ball expanding, $f$ has the shadowing property and the h-shadowing property. By these conditions, we conclude that $f$ is locally eventually onto, completing the proof of the theorem.
\end{proof}

This paper consists of two sections and two appendices. In the next section, we give several examples related to the main results. 

\section{Examples}

In this section, we present several examples concerning ball expanding maps and the main results in Section 1.

\begin{ex}
\normalfont
Let $X=\{0\}\cup\{2^{-n}\colon n\ge0\}$ and note that $X\subset[0,1]$. We define a surjective continuous map $f\colon X\to X$ by
\begin{itemize}
\item $f(1)=1$,
\item $f(x)=2x$ for all $x\in X\setminus\{1\}$.
\end{itemize}

We shall prove that $f$ is ball expanding.

\begin{proof}
Note that $y=f(2^{-1}y)$ for all $y\in X$. Let $0<\delta<2^{-1}$. It is sufficient to show that for any $x\in X$ and $y\in B_\delta(f(x))\setminus\{f(x)\}$, we have $y=f(z)$ for some $z\in B_{2^{-1}\delta}(x)$. If $x=1$, then $B_\delta(f(x))=B_\delta(1)=\{1\}$ and so $B_\delta(f(x))\setminus\{f(x)\}=\emptyset$. If $x\ne1$, then we have $f(x)=2x$. Letting $z=2^{-1}y$, we have $y=f(z)$. By
\[
|f(x)-y|=|2x-y|\le\delta,
\]
we obtain
\[
|x-z|=|x-2^{-1}y|\le2^{-1}\delta.
\]
This completes the proof of the claim.
\end{proof}

Note that $CR(f)=\{0,1\}$. Let $X^\mathbb{N}$ be the product space of infinitely many copies of $X$ and let $D$ denote the metric on $X^\mathbb{N}$ defined by
\[
D(x,y)=\sup_{j\ge1}2^{-j}|x_j-y_j|
\]
for all $x=(x_j)_{j\ge1},y=(y_j)_{j\ge1}\in X^\mathbb{N}$. We define $g\colon X^\mathbb{N}\to X^\mathbb{N}$ by: for any $x=(x_j)_{j\ge1},y=(y_j)_{j\ge1}\in X^\mathbb{N}$, $y=g(x)$ if and only if $y_j=f(x_j)$ for all $j\ge1$. Since $f$ is ball expanding, $f$ has the h-shadowing property and so does $g$. Note that $g$ satisfies the following properties
\begin{itemize}
\item $g(x)=x$ for all $x\in\{0,1\}^\mathbb{N}$,
\item $CR(g)=\{0,1\}^\mathbb{N}$, an uncountable set,
\item $\mathcal{C}(g)=\{\{x\}\colon x\in\{0,1\}^\mathbb{N}\}$ and $\mathcal{C}_{\rm ter}(g)=\{\{1^\infty\}\}$.
\end{itemize}
If $g$ is ball expanding, then since $h_{\rm top}(g)=0$, by Theorem 1.2, $CR(g)$ must be a finite set. It follows that $g$ is not ball expanding. Defining $x=(x_j)_{j\ge1}\in X^\mathbb{N}$ by $x_j=2^{-j+1}$ for all $j\ge1$, we easily see that
\[
\lim_{i\to\infty}D(g^i(x),1^\infty)=0;
\]
however,
\[
x\not\in\bigcup_{i\ge0}g^{-i}(CR(g))
\] 
(see Remark 1.6 in Section 1).
\end{ex}

\begin{ex}
\normalfont
Let $I_n=[4^{-n},2\cdot 4^{-n}]$ for all $n\ge1$. Let
\[
X=\{0,2\}\cup\bigcup_{n\ge1}I_n
\]
and note that $X\subset[0,2]$. We define a surjective continuous map $f\colon X\to X$ by
\begin{itemize}
\item $f(x)=2$ for all $x\in\{2\}\cup I_1$,
\item $f(x)=4x$ for all $x\in X\setminus(\{2\}\cup I_1)$.
\end{itemize}

We shall prove that $f$ is ball expanding.

\begin{proof}
Note that $y=f(4^{-1}y)$ for all $y\in X$. Let $0<\delta<3\cdot2^{-1}$. It is sufficient to show that for any $x\in X$ and $y\in B_\delta(f(x))\setminus\{f(x)\}$, we have $y=f(z)$ for some $z\in B_{4^{-1}\delta}(x)$. If $x\in\{2\}\cup I_1$, then $B_\delta(f(x))=B_\delta(2)=\{2\}$ and so $B_\delta(f(x))\setminus\{f(x)\}=\emptyset$. If $x\not\in\{2\}\cup I_1$, then we have $f(x)=4x$. Letting $z=4^{-1}y$, we have $y=f(z)$. By
\[
|f(x)-y|=|4x-y|\le\delta,
\]
we obtain
\[
|x-z|=|x-4^{-1}y|\le4^{-1}\delta;
\]
thus the claim has been proved.
\end{proof}

Note that $X$ is uncountable and not totally disconnected, but $h_{\rm top}(f)=0$.
\end{ex}

\begin{ex}
\normalfont
In \cite{BGO}, the tent map $T\colon[0,1]\to [0,1]$, defined as $T(x)=1-|1-2x|$ for all $x\in[0,1]$, is given as an example of a ball expanding map (see Example 2.9 of \cite{BGO}). For the sake of completeness, we give a proof.

\begin{proof}
It is sufficient to show that 
\[
B_{2\delta}(T(x))\subset T(B_\delta(x))
\]
for all $x\in[0,1]$ and $0<\delta\le4^{-1}$. We may assume $x\in[0,2^{-1}]$. Note that $T(x)=2x$. If $x\in[0,\delta]$, then $T(B_\delta(x))=[0,2x+2\delta]$. Since
\[
B_{2\delta}(T(x))\subset[2x-2\delta,2x+2\delta]\cap[0,1]=[0,2x+2\delta],
\]
we obtain $B_{2\delta}(T(x))\subset T(B_\delta(x))$. If $x\in[\delta,2^{-1}-\delta]$, then $T(B_\delta(x))=[2x-2\delta,2x+2\delta]$. Since
\[
B_{2\delta}(T(x))\subset[2x-2\delta,2x+2\delta],
\]
we obtain $B_{2\delta}(T(x))\subset T(B_\delta(x))$. If $x\in[2^{-1}-\delta,2^{-1}]$, then $T(B_\delta(x))=[2x-2\delta,1]$. Since
\[
B_{2\delta}(T(x))\subset[2x-2\delta,2x+2\delta]\cap[0,1]=[2x-2\delta,1],
\]
we obtain $B_{2\delta}(T(x))\subset T(B_\delta(x))$. This completes the proof of the claim.
\end{proof}
Note that $[0,1]$ is connected and $T$ is locally eventually onto.

Let $S^1=\{z\in\mathbb{C}\colon|z|=1\}$ with the arc-length metric. It is easy to see that the doubling map $f\colon S^1\to S^1$, defined as $f(z)=z^2$ for all $z\in S^1$, is ball expanding, where $S^1$ is connected and $f$ is locally eventually onto.

Let $g\colon[0,1]\to[0,1]$ be the logistic map defined as $g(x)=4x(1-x)$ for all $x\in[0,1]$. It is well-known that $g\colon[0,1]\to[0,1]$ is topologically conjugate to $T\colon[0,1]\to [0,1]$. As claimed in \cite{BGO}, since $T$ is ball expanding and so has the h-shadowing property, $g$ also has the h-shadowing property. Let $0<\delta\le2^{-1}$ and note that $g(2^{-1}-\delta)=g(2^{-1}+\delta)=1-4\delta^2$. We have
\[
B_\delta(g(2^{-1}))=B_\delta(1)=[1-\delta,1]
\]
and
\[
g(B_\delta(2^{-1}))=g([2^{-1}-\delta,2^{-1}+\delta])=[1-4\delta^2,1].
\]
If $\delta<4^{-1}$, then $1-\delta<1-4\delta^2$ and so $B_\delta(g(2^{-1}))\not\subset g(B_\delta(2^{-1}))$. This implies that $g$ is not ball expanding (see Example 5.5 of \cite{BGO}).
\end{ex}

\begin{ex}
\normalfont
Let $X^\mathbb{N}$ be the product space of infinitely many copies of $X$ and let $D$ denote the metric on $X^\mathbb{N}$ defined by
\[
D(x,y)=\sup_{j\ge1}2^{-j}d(x_j,y_j)
\]
for all $x=(x_j)_{j\ge1},y=(y_j)_{j\ge1}\in X^\mathbb{N}$. We consider the shift map $f\colon X^\mathbb{N}\to X^\mathbb{N}$ defined as: for any $x=(x_j)_{j\ge1},y=(y_j)_{j\ge1}\in X^\mathbb{N}$, $y=f(x)$ if and only if $y_j=x_{j+1}$ for all $j\ge1$.

We shall prove that $f$ is ball expanding.

\begin{proof}
It is sufficient to show that 
\[
B_\delta(f(x))\subset f(B_{2^{-1}\delta}(x))
\]
for all $x\in X^\mathbb{N}$ and $\delta>0$. Given any $x\in X^\mathbb{N}$ and $y\in B_\delta(f(x))$, we have $d(x_{j+1},y_j)\le2^j\delta$ for all $j\ge1$. We define $z=(z_j)_{j\ge1}\in X^\mathbb{N}$ by
\begin{itemize}
\item $z_1=x_1$,
\item $z_{j+1}=y_j$ for all $j\ge1$.
\end{itemize}
Note that $y=f(z)$. Letting $m_j=2^{-j-1}d(x_{j+1},z_{j+1})$, $j\ge0$, we see that
 \begin{itemize}
\item $m_0=2^{-1}d(x_1,x_1)=0$,
\item $m_j=2^{-j-1}d(x_{j+1},y_j)\le2^{-j-1}\cdot2^j\delta\le2^{-1}\delta$ for all $j\ge1$.
\end{itemize}
It follows that $D(x,z)=\sup_{j\ge0}m_j\le2^{-1}\delta$, i.e., $z\in B_{2^{-1}\delta}(x)$. Since $y\in B_\delta(f(x))$ is arbitrary, we conclude that
\[
B_\delta(f(x))\subset f(B_{2^{-1}\delta}(x)).
\]
Since $x\in X$ is arbitrary, this completes the proof of the claim.
\end{proof}

Note that $f\colon X^\mathbb{N}\to X^\mathbb{N}$ is locally eventually onto. If $X=\{0,1\}$, then $X^\mathbb{N}$ is perfect and totally disconnected. If $X=[0,1]$, then $X^\mathbb{N}$ is connected and infinite-dimensional.
\end{ex}

\appendix

\section{}

The aim of this Appendix A is to show that the same conclusion as in Lemma 1.1 holds under the assumption of $L$-Lipschitz shadowing property, where $0<L<1$. Precisely, we prove the following. 

\begin{lem}
Let $f\colon X\to X$ be a continuous map. Let $0<L<1$ and $\delta_0>0$ be as in the definition of $L$-Lipschitz shadowing property. For any $C\in\mathcal{C}(f)$ and $x\in X$, if $d(x,C)\le\delta_0$ and
\[
\lim_{i\to\infty}d(f^i(x),C)=0,
\]
then $x\in C$.
\end{lem}

\begin{proof}
Let us show that for every $i\ge0$, we have
\[
\max\{d(x_i,C),d(f^i(x_i),x)\}\le L^i\delta_0
\]
for some $x_i\in X$. We argue by induction on $i$. When $i=0$, let $x_0=x$. Assume that the inequality holds for some $i\ge0$ and $x_i\in X$. We take $y,z\in C$ such that
\[
d(x_i,y)=d(x_i,C)\le L^i\delta_0
\]
and $y=f(z)$. Since
\[
\gamma_i=(z,x_i,f(x_i),\dots,f^{i-1}(x_i),x)
\]
($\gamma_0=(z,x)$ if $i=0$) is an $L^i\delta_0$-chain of $f$, we obtain
\[
\max\{d(x_{i+1},z),d(f^{i+1}(x_{i+1}),x)\}\le L^{i+1}\delta_0
\]
for some $x_{i+1}\in X$. Since
\[
d(x_{i+1},C)\le d(x_{i+1},z)\le L^{i+1}\delta_0,
\]
the induction is complete. From
\[
\lim_{i\to\infty}d(x_i,C)=\lim_{i\to\infty}d(f^i(x_i),x)=0,
\]
it follows that $w\rightarrow x$ for some $w\in C$. The rest of the proof is identical to that of Lemma 1.1.
\end{proof}

\section{}

By Theorem 4.3 of \cite{BGO}, we know that if a continuous map $f\colon X\to X$ is ball expanding, then $f$ has the h-shadowing property. The aim of this Appendix B is to prove the following generalization of Theorem 1.5.

\begin{thm}
If $X$ is connected and if a continuous map $f\colon X\to X$ has the h-shadowing property, then $f$ is locally eventually onto. 
\end{thm}

For a subset $S$ of $X$ and $r>0$, we denote by $B_r(S)$ the closed $r$-neighborhood of $S$:
\[
B_r(S)=\{x\in X\colon d(x,S)=\inf_{y\in S}d(x,y)\le r\}.
\]
Let us first prove the following lemma.

\begin{lem}
Let $f\colon X\to X$ be a continuous map. For a closed subset $S$ of $X$, if
\begin{itemize}
\item $f(S)=S$,
\item $S$ is chain stable,
\end{itemize}
then for any $r>0$, there is a subset $A$ of $X$ such that $S\subset A\subset B_r(S)$ and $f(\overline{A})\subset{\rm int}[A]$, where ${\rm int}[A]$ denotes the interior of $A$.
\end{lem}

\begin{proof}
For $\delta>0$, let $A$ be the set of $y\in X$ such that there are $x\in S$ and a $\delta$-chain $(x_i)_{i=0}^k$ of $f$ such that $x_0=x$ and $x_k=y$. For any $y\in S$, since $f(S)=S$, by taking $x\in S$ with $y=f(x)$, we have that $(x,y)$ is a $\delta$-chain of $f$, thus $y\in A$. It follows that $S\subset A$. Since $S$ is chain stable, for any $r>0$, we have $A\subset B_r(S)$ for some $\delta>0$.

For any $z\in\overline{A}$, we take $b>0$ such that $f(B_b(z))\subset B_{\delta/2}(f(z))$. Since $z\in\overline{A}$, we have $y\in B_b(z)$ for some $y\in A$. Since $y\in A$, there are $x\in S$ and a $\delta$-chain $(x_i)_{i=0}^k$ of $f$ such that $x_0=x$ and $x_k=y$. For any $w\in B_{\delta/2}(f(z))$, by
\[
d(f(y),w)\le d(f(y),f(z))+d(f(z),w)\le\delta/2+\delta/2=\delta,
\]
we see that
\[
\gamma=(x_0,x_1,\dots,x_k,w)
\]
is a $\delta$-chain of $f$; therefore, $w\in A$. It follows that $B_{\delta/2}(f(z))\subset A$, thus $f(z)\in{\rm int}[A]$. Since $z\in\overline{A}$ is arbitrary, we obtain $f(\overline{A})\subset{\rm int}[A]$, completing the proof of the lemma.
\end{proof}

With the aid of Lemma B.1, we prove the following lemma.

\begin{lem}
Let $f\colon X\to X$ be a continuous map with the h-shadowing property. For a closed subset $S$ of $X$, if
\begin{itemize}
\item $f(S)=S$,
\item $S$ is chain stable,
\end{itemize}
then for any $r>0$, there is a clopen subset $B$ of $X$ such that $S\subset B\subset B_r(S)$.
\end{lem}

\begin{proof}
For $r>0$, we take a subset $A$ of $X$ as given in Lemma B.1. Note that $f(\overline{A})\subset\overline{A}$ and let $B=\bigcap_{i\ge0}f^i(\overline{A})$. We have the following properties:
\begin{itemize}
\item $S\subset B\subset{\rm int}[A]\subset  B_r(S)$,
\item $B$ is closed in $X$ and satisfies $f(B)=B$.
\end{itemize}
Let us prove that $B$ is open (and so clopen) in $X$. We take $\epsilon>0$ such that $B_\epsilon(B)\subset A$. For any $a>0$, by compactness, there is $i\ge0$ such that $f^j(\overline{A})\subset B_a(B)$ for all $j\ge i$. It follows that $f^j(B_\epsilon (B))\subset B_a(B)$ for all $j\ge i$. For this $\epsilon$, we fix $\delta>0$ as in the definition of the h-shadowing property. If $B$ is not open in $X$, then we have
\[
0<d(x,B)\le\delta
\]
for some $x\in X$. We take $a>0$ and $j\ge1$ with $a<d(x,B)$ and  $f^j(B_\epsilon(B))\subset B_a(B)$. There are $y,z\in B$ such that 
\[
d(x,y)=d(x,B)\le\delta
\]
and $y=f^j(z)$. Since
\[
\gamma=(z,f(z),\dots,f^{j-1}(z),x)
\]
is a $\delta$-chain of $f$, we obtain $x=f^j(w)$ for some $w\in B_\epsilon(z)$. It follows that
\[
x\in f^j(B_\epsilon(B))\subset B_a(B),
\]
i.e., $d(x,B)\le a$, a contradiction. Thus, we conclude that $B$ is open in $X$, completing the proof of the lemma.
\end{proof}

Given a continuous map $f\colon X\to X$ and $x\in X$, we define the {\em chain $\omega$-limit set} $\omega^\ast(x,f)$ of $x$ for $f$ as the set of $y\in X$ such that for any $\delta>0$ and $l>0$, there is a $\delta$-chain $(x_i)_{i=0}^k$ of $f$ with $x_0=x$, $x_k=y$, and $k\ge l$. Note that for every $x\in X$,
\begin{itemize}
\item $\omega^\ast(x,f)$ is closed in $X$,
\item $f(\omega^\ast(x,f))=\omega^\ast(x,f)$,
\item $\omega^\ast(x,f)$ is chain stable, because $y\rightarrow z$ implies $z\in\omega^\ast(x,f)$ for all $y\in\omega^\ast(x,f)$ and $z\in X$.
\end{itemize}

Finally, we prove Theorem B.1.

\begin{proof}[Proof of Theorem B.1]
Let $x\in X$. Since $f$ has the h-shadowing property, Lemma B.2 implies that for any $r>0$, there is a clopen subset $B$ of $X$ such that
\[
\omega^\ast(x,f)\subset B\subset B_r(\omega^\ast(x,f)).
\]
As $X$ is connected, we obtain $B=X$ and so $X=B_r(\omega^\ast(x,f))$. Since $r>0$ is arbitrary, we obtain $X=\omega^\ast(x,f)$. Since $x\in X$ is arbitrary, by the definition of the chain $\omega$-limit set, we obtain $x\rightarrow y$ for all $x,y\in X$; therefore, $\mathcal{C}(f)=\{X\}$. By Corollary 14 of \cite{RW}, $f$ is chain mixing. Since $f$ has the h-shadowing property, we conclude that $f$ is locally eventually onto, completing the proof of the theorem.
\end{proof}

\end{document}